# Ergotropic advantage in a measurement-fueled quantum heat engine

### v1

@@ Abstract
This paper investigates a coupled two-qubits heat engine fueled by generalized measurements of the spin components and using a single heat reservoir as sink. Our model extends the four-stroke engine proposed by Yi and coworkers [Phys. Rev. E 96, 022108 (2017)] by introducing an ergotropy-extracting stroke, resulting in a five-stroke cycle. For measurements along z-z directions, we find two possible occupation distributions that yield an active state and the ergotropic stroke improves the performance of the engine over the four-stroke cycle. Further, the three-stroke engine (without the adiabatic strokes) yields the same performance as the five-stroke engine. For arbitrary working medium and non-selective measurements, we prove that the total work output of a five-stroke engine is equal to the sum of the work outputs of the corresponding four-stroke and three-stroke engines. For measurement directions other than z-z, there may be many possible orderings of the post-measurement probabilities that yield an active state. However, as we illustrate for specific cases (x-x and x-z directions), a definite ordering is obtained with the projective measurements. Thus, we find that the five-stroke engine exploiting ergotropy outperforms both its four-stroke as well as three-stroke counterparts.
@@ I Introduction
Quantum thermodynamics [1, 2, 3, 4, 5, 6, 7], involves, amongst other pursuits, the study of thermal machines based on a quantum working medium. The idea of a quantum heat engine was first proposed by Scovil and Schulz-DuBois [8], who argued that a three-level maser, in simultaneous contact with a hot and a cold reservoir, can be regarded as a heat engine. State-of-the-art technology now allows us to experimentally control systems at the quantum scale, such as spins [9, 10, 11, 12, 13, 14], superconducting qubits [15, 16] and trapped ions [17, 18]. So, the non-classical resources such as quantum entanglement, quantum coherence, quantum correlations and noise are being hotly pursued for their use in improving the performance of thermal machines away from the classical regime [19, 20, 21, 22, 23, 24, 25, 26, 27, 28].
@@
In recent years, quantum measurements [29, 30] have attracted increasing attention due to their ability to mediate energy exchange—either as work extraction [31] or as effective heat transfer. Measurement-induced energy injection can mimic the role of a hot reservoir, provided the measured observable does not commute with the system Hamiltonian [32, 33, 34, 35]. Importantly, such engines operate without feedback control, distinguishing them from Szilard-type engines [36, 37, 38]. Related developments include quantum refrigerators and cooling protocols [39, 40], weak and continuous measurement schemes [41, 42, 43], non-ideal measurements [44] and work-fluctuation analyses [45]. Measurements further induce entropy production, probe–system correlations, and raise questions about associated energetic costs [46, 47]. The classification of measurement-induced energy exchange and its role as a fueling mechanism [48] remains an active research direction.
@@
The present work extends the measurement-fueled engine [32, 33, 49] by using weak quantum measurements on a coupled-qubits system while exploiting the active nature of the post-measurement state of the system. We incorporate an ergotropy-extracting stroke [50, 51, 52, 53, 54, 55] after the measurement stroke, thereby upgrading the four-stroke cycle into a five-stroke cycle. With measurements on the spin components along specific directions, we highlight the thermodynamic advantage of the ergotropic stroke by comparing the performance of five-stroke engine with the four-stroke (without ergotropy) as well as a three-stroke (without adiabatic strokes) cycle.
@@
The paper is organized as follows. Section II describes the five-stroke heat cycle of the measurement-fueled engine, which is applied to a coupled qubits working medium in Section III.1 for z-z measurements, while comparing with four-stroke and three-stroke cycle in Section III.2 and III.3 respectively. Section IV describes the results for other measurement directions. Section V concludes the work with a summary and outlook.
@@ II Five-stroke engine
The working medium, or the system, is described by the Hamiltonian $H(B)$, with its initial state as the thermal state ($\rho^{\rm in}$) corresponding to a heat reservoir at inverse temperature $\beta$, and the control parameter set at $B=B_{2}$. The occupation probability for the $n$-th energy eigenstate is labelled as $p_{n}$. The system undergoes a heat cycle consisting of five strokes, as described below (see also Fig. 1).
Figure 1: A five-stroke engine cycle. 1 $\rightarrow$ 2 is the first adiabatic stroke. 2 $\rightarrow$ 3 is the measurement stroke. 3 $\rightarrow$ 4 is the ergotropy extraction stroke. 4 $\rightarrow$ 5 is the second adiabatic stroke. 5 $\rightarrow$ 1 is the thermalization stroke.
Stroke $1\to 2$: This is a quantum adiabatic process (AP-1) in which the parameter $B$ is increased from $B_{2}$ to $B_{1}$ sufficiently slowly such that the quantum adiabatic theorem [56] holds and the occupation probabilities remain unchanged.
@@
Stroke $3\to 4$: This is termed as the ergotropic stroke which is executed only if the post-measurement state is an active state [50]. In that case, ergotropy is the maximum work extracted from the system via a cyclic unitary evolution of its Hamiltonian.
@@
Suppose that after ergotropy extraction ($3\to 4$), the system is left in a passive state $\rho^{\prime}$ with occupation probabilities $p_{n}^{\prime}$. Then, the ergotropy is given by
@@
Finally, the heat exchanged with the reservoir is
$Q_{\rm res}=\sum_{n}E_{n}(B_{2})[p_{n}-p^{\prime}_{n}].$ (5)
@@ III Coupled-qubits working medium
We consider two coupled qubits following a 1D Hamiltonian with isotropic Heisenberg interaction:
@@
In the first adiabatic stroke, the work [Eq. (1)] is evaluated to be
@@
Note that $W_{1}$ is independent of the measurement choice as it depends only on the initial state and the Hamiltonian $H(B)$.
@@
$\begin{split}M_{\pm,\pm}&=(c_{0}I\pm c_{1}\vec{\sigma}^{\rm A}.\hat{n}^{\rm A})\otimes(c_{0}I\pm c_{1}\vec{\sigma}^{\rm B}.\hat{n}^{\rm B}),\\ M_{\pm,\mp}&=(c_{0}I\pm c_{1}\vec{\sigma}^{\rm A}.\hat{n}^{\rm A})\otimes(c_{0}I\mp c_{1}\vec{\sigma}^{\rm B}.\hat{n}^{\rm B}),\end{split}$ (10)
where $c_{0}$ and $c_{1}$ are real parameters. From the completeness relation: $2c_{0}^{2}+2c_{1}^{2}=1$. One of the parameters, say $c_{0}$, can be chosen to define the strength of the measurement. Thus, $c_{0}=c_{1}=1/2$ implies a strong or projective measurement [33]. The general expressions for the post-measurement probabilities are quite complicated and it is not feasible to determine the conditions for their relative ordering. So, in the following, we study measurements along specific directions.
@@
We choose $\vec{\sigma}^{\rm A}.\hat{n}^{\rm A}=\sigma_{z}^{\rm A}$ and $\vec{\sigma}^{\rm B}.\hat{n}^{\rm B}=\sigma_{z}^{\rm B}$. The post-measurement probabilities are given by
@@
As a result of the measurement, the occupation probabilities of levels $E_{2}$ and $E_{4}$ are unchanged, while those of $E_{1}$ and $E_{3}$ come closer to each other. Intuitively, we see that the post-measurement probability distribution becomes more uniform and so its Shannon entropy increases. Also, $p^{\rm{PM}}_{1}-p^{\rm{PM}}_{3}=(4c_{0}^{2}-1)^{2}(p_{1}-p_{3})\geq 0$, where the equality is obtained for $c_{0}=1/2$. Note that $p_{4}$ is still the lowest probability in the distribution (11) since the probabilities were initially ordered as $p_{1}>p_{2}>p_{3}>p_{4}$. Based on these considerations, we can infer: $p^{\rm{PM}}_{1}\geq p^{\rm{PM}}_{3}>p_{4}$. However, to ascertain the active nature of the state $\rho^{\rm{PM}}$, we have to determine the relative ordering of $p_{2}$ also. It can be seen that an active post-measurement state implies one of the following two possibilities:
@@
[…]
@@
It is interesting to note that $\eta$ is independent of the parameter $B_{1}$. This is not a generic feature, but depends on the energy spectrum. For the given model, $\eta$ is maximum at $c_{0}=1/2$ i.e. for projective measurements. The efficiency is plotted in Fig. 3. For a given $c_{0}$, $\eta$ increases with $\beta$. For low enough temperatures, $\eta$ grows rapidly with $c_{0}$ and approaches its limiting value $B_{2}/4J$. In Fig. 4, the efficiency is plotted versus $J$ for a fixed $c_{0}$ at various temperatures. Again, decreasing the temperature increases the efficiency at low $J$ values, while for large $J$ values, all efficiency curves merge into each other and goes to zero at very large $J$.
@@
Figure 3: R1 case: Efficiency [Eq. (19)] vs. $c_{0}\in[0,1/\sqrt{2}]$. Red curve ($\beta=1$), green curve ($\beta=2$), blue curve ($\beta=3$) and magenta curve ($\beta=4$). Here, $B_{2}=3$ and $J=1$. So, the limiting value of the efficiency (as $\beta\to\infty$) is $B_{2}/4J=0.75$.
Figure 4: R1 case: Efficiency [Eq. (19)] vs. $J$. Red curve ($\beta=1$), green curve ($\beta=2$), blue curve ($\beta=3$) and magenta curve ($\beta=4$). Here, $B_{2}=3$ and $c_{0}=0.3$.
@@ III.2 Four-stroke cycle
For comparison, we also consider the four-stroke cycle [32, 33] that does not involve the ergotropic stroke. It implies the occupation probabilities in AP-2 are given by $p^{\rm{PM}}_{n}$. The expressions for heat and work in different stages can be obtained by replacing each $p_{n}^{\prime}$ with $p^{{\rm PM}}_{n}$ in the expressions for the five-stroke cycle as in Sec. II. Thus, the total work extracted can be expressed as:
@@
For the two-qubits system and with z-z measurements, we obtain So, the net work output of the four-stroke cycle is zero. This feature was noted in Ref. [33] for z-z projective measurements and we affirm it for generalized measurements too. Thus, we can conclude that the ergotropy stroke improves the performance of the engine based on generalized z-z measurements.
Figure 5: R2 case: Efficiency vs. $c_{0}$. Red curve ($\beta=1$), green curve ($\beta=2$), blue curve ($\beta=3$) and magenta curve ($\beta=4$). $B_{2}=3.9$, $J=1$. The efficiency decreases as the temperature is lowered at a given measurement strength, unlike the R1 case (see Fig. 3). The R2 efficiency is also maximized at $c_{0}=1/2$.
@@ III.3 Three-stroke cycle
As an alternative, we consider an ergotropy based three-stroke cycle consisting of the following steps. We prepare the system with Hamiltonian $H(B_{2})$ in a thermal state $\rho^{\rm in}$ at inverse temperature $\beta$. i) The system is isolated from the reservoir (assuming a weak interaction with the reservoir and so negligible costs in attaching and detaching the system from the reservoir) and nonselective generalized measurements are performed which transform the state to $\rho^{\rm PM}$. ii) Ergotropy, $W_{\rm erg}$, is extracted which yields the passive state $\rho^{\prime}$. iii) Finally, the system is brought to the initial state by thermalization with the reservoir. The net work extracted, equal in magnitude to the ergotropy in step (ii), is given by
@@
For the R1 case, we obtain $W_{\rm T}^{(3)}=2B_{2}(p^{\rm{PM}}_{3}-p_{2})$, which, as expected, is equal to the work in the five-stroke cycle [Eq. (18)] since $W_{\rm T}^{(4)}=0$ for z-z measurements. $Q_{\rm M}$ for the three-stroke cycle is also given by Eq. (12). Therefore, its efficiency is the same as for the five-stroke cycle [Eq. (19)]. A similar conclusion holds for the case R2.

### v2

@@ Abstract
This paper investigates a coupled two-qubits heat engine fueled by generalized measurements of the spin components and using a single heat reservoir as sink. Our model extends the proposal of Yi and coworkers [Phys. Rev. E 96, 022108 (2017)] where the role of a hot reservoir in a four-stroke cycle was replaced by a quantum measurement apparatus, the other steps being two quantum adiabatic strokes and thermalization with a cold reservoir. We propose a five-stroke cycle, where an ergotropy extracting stroke is introduced following the measurement stroke, and study the effect of measurements of different spin components on the performance of the machine. For measurements along z-z directions, we find two possible occupation distributions that yield an active state and the ergotropic stroke improves the performance of the engine over the four-stroke cycle. Further, the three-stroke engine ( without the adiabatic strokes) yields the same performance as the five-stroke engine. For arbitrary working medium and non-selective measurements, we prove that the total work output of a five-stroke engine is equal to the sum of the work outputs of the corresponding four-stroke and three-stroke engines. For measurement directions other than z-z, there may be many possible orderings of the post-measurement probabilities that yield an active state. However, as we illustrate, for specific cases (e.g. x-x), a definite ordering may be obtained with the projective measurements. Thus, we find that the five-stroke engine exploiting ergotropy outperforms both its four-stroke as well as three-stroke counterparts.
@@ I Introduction
Quantum thermodynamics [1, 2, 3, 4, 5, 6, 7] involves, amongst other pursuits, the study of thermal machines based on a quantum working medium. The idea of a quantum heat engine was first proposed by Scovil and Schulz-DuBois [8], who argued that a three-level maser, in simultaneous contact with a hot and a cold reservoir, can be regarded as a heat engine. State-of-the-art technology now allows us to experimentally control systems at the quantum scale, such as spins [9, 10, 11, 12, 13, 14, 15], superconducting qubits [16, 17] and trapped ions [18, 19]. So, the non-classical resources such as entanglement, quantum coherence, quantum correlations and noise are being hotly pursued for their use in improving the performance of thermal machines away from the classical regime [20, 21, 22, 23, 24, 25, 26, 27, 28, 29].
@@
In recent years, quantum measurements [30, 31] have attracted increasing attention due to their ability to mediate energy exchange—either as work extraction [32] or as effective heat transfer. Measurement-induced energy injection can mimic the role of a hot reservoir, provided the measured observable does not commute with the system Hamiltonian [33, 34, 35, 36]. Importantly, such engines operate without feedback control, distinguishing them from Szilard-type engines [37, 38, 39]. Related developments include quantum refrigerators and cooling protocols [40, 41], weak and continuous measurement schemes [42, 43, 44], non-ideal measurements [45] and work-fluctuation analyses [46]. Measurements further induce entropy production, probe–system correlations, and raise questions about associated energetic costs [47, 48, 49]. The classification of measurement-induced energy exchange and its role as a fueling mechanism [50, 51] remains an active research direction.
@@
The main idea of our investigation is that after the non-selective quantum measurement injects heat into the working medium, the latter may be left in a non-passive or active state. From this state, it is possible to extract work by a unitary process in which the Hamiltonian of the system undergoes a cyclic evolution and the maximum work so extracted is known as the ergotropy [52]. The present work extends the measurement-fueled engine [33, 34, 53] by using weak quantum measurements on a coupled-qubits system while exploiting the active nature of the post-measurement state of the system. We incorporate an ergotropy-extracting stroke [54, 55, 56, 57, 58] after the measurement stroke, thereby upgrading the four-stroke cycle to a five-stroke cycle. With measurements on the spin components along specific directions, we highlight the thermodynamic advantage of the ergotropic stroke by comparing the performance of the five-stroke engine with the four-stroke (without ergotropy) as well as a three-stroke (without adiabatic strokes) cycle.
@@
The paper is organized as follows. Section II describes the five-stroke heat cycle of the measurement-fueled engine, which is applied to a coupled qubits working medium with z-z measurements in Section III.1, while comparing with the four-stroke and the three-stroke cycles in Section III.2 and III.3 respectively. Section IV describes the results for other measurement directions. We conclude our analysis in Section V.
@@ II Five-stroke engine
The working medium, or the system, is described by the Hamiltonian $H(B)$, with its initial state as the thermal state ($\rho^{\rm in}$) corresponding to a heat reservoir at inverse temperature $\beta$, and the control parameter set at $B=B_{2}$. We follow the convention of the two-reservoir model in which the (magnetic field) parameter at the cold reservoir is labelled as $B_{2}$ and in the present case, the only available reservoir acts as a cold reservoir. The occupation probability for the $n^{\rm th}$ energy eigenstate is labelled as $p_{n}$. The system undergoes a heat cycle consisting of five strokes, as described below (see Fig. 1).
Figure 1: A five-stroke engine cycle consisting of the first quantum adiabatic stroke ($1\rightarrow 2$), the measurement stroke ($2\rightarrow 3$), the ergotropy extraction stroke ($3\rightarrow 4$), the second quantum adiabatic stroke ($4\rightarrow 5$) and finally the thermalization stroke with the heat reservoir ($5\rightarrow 1$). Without the stroke $3\rightarrow 4$, the cycle reduces to the four-stroke engine of Ref. [33].
Stroke $1\to 2$: This is a quantum adiabatic process (AP-1) in which the parameter $B$ is increased from $B_{2}$ to $B_{1}$ sufficiently slowly such that the quantum adiabatic theorem [59] holds and the occupation probabilities remain unchanged. Since the system remains in its instantaneous energy eigenstate, the state at the end of this stroke is given by: $\rho=\sum_{n}p_{n}|n,B_{1}\rangle\langle n,B_{1}|$, where $|n,B_{1}\rangle$ are energy eigenstates at the end of this process.
@@
Stroke $3\to 4$: This is termed as the ergotropic stroke. Following Ref. [52], the ergotropy is defined as the maximum work extractable from a quantum system via a cyclic unitary on its Hamiltonian. The state of the system from which work can be so extracted is denoted as an active state, otherwise it is called a passive state. Thus, the stroke $3\to 4$ is executed only if the post-measurement state is an active state (the criterion is explained in the next Section). In other words, if the post-measurement state is a passive state, then this stroke is not executed and the proposed heat cycle reduces to the four-stroke cycle of Ref. [33], with all the other steps remaining the same.
@@
To implement the ergotropic stroke, suppose that under the action of a cyclic unitary $U$, the post-measurement state $\rho^{\rm{PM}}$ transforms to $U\rho^{\rm PM}U^{{\dagger}}$ while the Hamiltonian returns to its original form $H(B_{1})$. The average work extracted in this process is given by $W(\rho^{\rm{PM}},U)={\rm Tr}[H(U\rho^{\rm{PM}}U^{{\dagger}}-\rho^{\rm{PM}})]<0$. The maximum of $W(\rho^{\rm{PM}},U)$ over all cyclic unitaries is called the ergotropy $W_{\rm erg}$. After the extraction of ergotropy (stroke $3\to 4$), the passive state is given by $|\rho^{\prime}=\sum_{n}p_{n}^{\prime}\ket{n,B_{1}}\bra{n,B_{1}}$, where due to the unitary nature of the ergotropic stroke, $p_{n}^{\prime}$ are just the shuffled eigenvalues of $\rho^{\rm PM}$ such that $p_{n}^{\prime}>p_{n+1}^{\prime}$ for $E_{n}(B_{1})<E_{n+1}(B_{1})$. Then, the ergotropy is given by
@@
Finally, the heat rejected to the reservoir is
$Q_{\rm res}=\sum_{n}E_{n}(B_{2})[p_{n}-p^{\prime}_{n}]<0.$ (5)
@@ III Coupled-qubits working medium
We consider two coupled qubits following a 1D Hamiltonian with isotropic Heisenberg interaction [60]:
@@
Now, the initial state is $\rho^{\rm{in}}={e^{-\beta H(B_{2})}}/{Z}$, where $Z=\sum_{n=1}^{4}e^{-\beta E_{n}}$ is the partition function. Then, we have $p_{n}=e^{-\beta E_{n}}/Z$. In the first adiabatic stroke, the work [Eq. (1)] is evaluated to be
@@
$\begin{split}M_{\pm,\pm}&=(c_{0}I^{\rm A}\pm c_{1}\vec{\sigma}^{\rm A}.\hat{n}^{\rm A})\otimes(c_{0}I^{\rm B}\pm c_{1}\vec{\sigma}^{\rm B}.\hat{n}^{\rm B})\\ M_{\pm,\mp}&=(c_{0}I^{\rm A}\pm c_{1}\vec{\sigma}^{\rm A}.\hat{n}^{\rm A})\otimes(c_{0}I^{\rm B}\mp c_{1}\vec{\sigma}^{\rm B}.\hat{n}^{\rm B})\end{split}$ (10)
where $c_{0}$ and $c_{1}$ are real parameters. From the completeness relation: $2c_{0}^{2}+2c_{1}^{2}=1$. One of the parameters, say $c_{0}$, can be chosen to define the strength of the measurement. Thus, $c_{0}=c_{1}=1/2$ implies a strong or projective measurement [34]. The general expressions for the post-measurement probabilities are quite complicated and it is not feasible to determine the conditions for their relative ordering. Now, we study measurements along specific directions.
@@
We choose $\vec{\sigma}^{\rm A}.\hat{n}^{\rm A}=\sigma_{z}^{\rm A}$ and $\vec{\sigma}^{\rm B}.\hat{n}^{\rm B}=\sigma_{z}^{\rm B}$. The post-measurement occupation probabilities are given by
@@
As a result of the measurement, the occupation probabilities of levels $E_{2}$ and $E_{4}$ are unchanged, while those of $E_{1}$ and $E_{3}$ come closer to each other. Intuitively, we see that the post-measurement probability distribution becomes more uniform and so its Shannon entropy increases. Also, $p^{\rm{PM}}_{1}-p^{\rm{PM}}_{3}=(4c_{0}^{2}-1)^{2}(p_{1}-p_{3})\geq 0$, where the equality is obtained for $c_{0}=1/2$. Note that $p_{4}$ is still the lowest probability in the distribution (11) since the probabilities were initially ordered as $p_{1}>p_{2}>p_{3}>p_{4}$. Based on these considerations, we can infer: $p^{\rm{PM}}_{1}\geq p^{\rm{PM}}_{3}>p_{4}$. Note that $\rho^{\rm PM}$ has no coherence in the energy eigenbasis for the z-z case. Therefore, to ascertain the active nature of the state $\rho^{\rm{PM}}$, we have to determine the relative ordering of the probabilities $p^{\rm{PM}}_{n}$. It can be seen that an active post-measurement state implies one of the following two possibilities:
@@
Figure 2: R1 case: Heat and work contributions in the five-stroke cycle vs. $c_{0}$. Parameters are set at $B_{1}=3.5$, $B_{2}$ = 3, $J$ = 1 and $\beta$ = 1.
@@
It is interesting to note that $\eta$ is independent of the parameter $B_{1}$. This is not a generic feature, but depends on the energy spectrum. For the given model, $\eta$ is maximum at $c_{0}=1/2$ i.e. for projective measurements. The efficiency is plotted in Fig. 3. For a given $c_{0}$, $\eta$ increases with $\beta$. For low enough temperatures, $\eta$ grows rapidly with $c_{0}$ and approaches its limiting value $B_{2}/4J$. In Fig. 4, the efficiency is plotted versus $J$ for a fixed $c_{0}$ at various temperatures. Again, decreasing the temperature increases the efficiency at low $J$ values, while for high $J$ values, all efficiency curves merge into each other and goes to zero.
@@
Figure 3: R1 case: Efficiency [Eq. (19)] vs. $c_{0}\in[0,1/\sqrt{2}]$. $B_{2}=3$ and $J=1$. So, the limiting value of the efficiency (as $\beta\to\infty$) is $B_{2}/4J=0.75$.
Figure 4: R1 case: Efficiency [Eq. (19)] vs. $J$. $B_{2}=3$ and $c_{0}=0.3$.
@@ III.2 Four-stroke cycle
For comparison, we also consider the four-stroke cycle [33, 34] that does not involve the ergotropic stroke. It implies the occupation probabilities in AP-2 are given by $p^{\rm{PM}}_{n}$. The expressions for $W_{2}$ and $Q_{\rm res}$ are obtained by replacing $p_{n}^{\prime}$ with $p^{{\rm PM}}_{n}$ in the corresponding expressions for the five-stroke cycle as in Sec. II. Thus, the total work extracted can be expressed as:
@@
For the two-qubits system and with z-z measurements, we obtain that the net work output of the four-stroke cycle is zero. This feature was noted in Ref. [34] for z-z projective measurements and we affirm it for generalized measurements too. Thus, we can conclude that the ergotropy stroke improves the performance of the engine based on generalized z-z measurements.
Figure 5: R2 case: Efficiency vs. $c_{0}$. $B_{2}=3.9$, $J=1$. The efficiency decreases as the temperature is lowered at a given measurement strength, unlike the R1 case (see Fig. 3). The R2 efficiency is also maximized at $c_{0}=1/2$.
@@ III.3 Three-stroke cycle
As an alternative, we consider an ergotropy based three-stroke cycle consisting of the following steps. We prepare the system with Hamiltonian $H(B_{2})$ in a thermal state $\rho^{\rm in}$ at inverse temperature $\beta$. i) The system is isolated from the reservoir (assuming a weak interaction with the reservoir and so negligible costs in attaching and detaching the system from the reservoir) and nonselective generalized measurements are performed which transform the state to $\rho^{\rm PM}$. Heat transferred due to measurement is given by
$Q_{\rm M}=\sum_{n}E_{n}(B_{2})[p^{\rm{\rm PM}}_{n}-p_{n}]>0.$ (22)
ii) Ergotropy, $W_{\rm erg}$, is extracted which yields the passive state $\rho^{\prime}$. iii) Finally, the system is brought to the initial state by thermalization with the reservoir. The net work extracted, equal in magnitude to the ergotropy in step (ii), is given by
@@
For the R1 case, we obtain $W_{\rm T}^{(3)}=2B_{2}(p^{\rm{PM}}_{3}-p_{2})$, which, as expected, is equal to the work in the five-stroke cycle [Eq. (18)] since $W_{\rm T}^{(4)}=0$ for z-z measurements. $Q_{\rm M}=8J(p^{\rm{PM}}_{3}-p_{3})>0$ for the three-stroke cycle, which is the same as Eq. (12). Therefore, its efficiency is the same as for the five-stroke cycle [Eq. (19)] in the case of z-z measurements. A similar conclusion holds for the case R2.